B22 Calculate the root-mean-square speeds of Hz 8 Os mdecules at 20°C.
Answer #1

RMS velocity= square root (3RT/M)

R=8.314 kgm2/sec 2 K.mol

T is absolute temperature=273+20=293 K

Mass is in kg: O2: 3.2x10^-2 kg.mol

For H2: 0.2x10-2 kg.mol

For O2: square t(3x8.314x293/3.2x10-2)=477.8 m/sec

For H2: square root (3x8.314x293/0.2x10^-2)= 1911. 5 m/sec

Units for RMS velocity are m/sec will get on cancellation.